# Copositive criteria for a two-component dark matter model

### v1

@@ 5.3 Relic density and direct detection constraints
Figure 10: Results of Case VIII with $\lambda_{13},\lambda_{23},\lambda_{24},\lambda_{14}\leqslant 0$.
We give the results of Case VI in Fig. 8 with $\lambda_{12},\lambda_{13},\lambda_{14}\leqslant 0$, where the left picture corresponds to the viable parameter space of $|\lambda_{13}|-|\lambda_{23}|$ and the right picture is $|\lambda_{14}|-|\lambda_{24}|$. The main feature of this case is that the upper bound for $(|\lambda_{14}|,|\lambda_{24}|)$ increases to about 0.6 due to the interference effect when considering the signs of the couplings.
We give the results of Case VII in Fig. 9 and Case VIII in Fig. 10, where Fig. 9(a) and Fig. 9(d) are the results of $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$, Fig. 9(b) and Fig. 9(e) are results of $\lambda_{14},\lambda_{24},\lambda_{12}\leqslant 0$, and Fig. 9(c) and Fig. 9(f) are results of $\lambda_{12},\lambda_{13},\lambda_{24}\leqslant 0$. According to Fig. 9(a), the larger $(|\lambda_{13}|,|\lambda_{23}|)$ are excluded by dark matter relic density constraint for $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$ while for the other two cases, the larger values of $|\lambda_{13}|$ and $|\lambda_{23}|$ can rid from the constraints due to the interference effect. Similar to the above-mentioned cases, the shapes of the parameter space of $(|\lambda_{14}|,|\lambda_{24}|)$ are slightly different for the three cases according to Fig. 9(d) to Fig. 9(f). For Case VIII, we have $\lambda_{13},\lambda_{23},\lambda_{24},\lambda_{14}\leqslant 0$, and Fig. 10(a) shows the viable parameter space of $(|\lambda_{13}|,|\lambda_{23}|)$ while Fig. 10(b) is the result of $(|\lambda_{14}|,|\lambda_{24}|)$.
In conclusion, according to Fig. 2 to Fig. 10, when we consider the signs of the couplings, we will have different viable parameter spaces under the different copositive criteria. We consider the 17 different cases and randomly scan the chosen parameter space under dark matter relic density constraint and direct detection constraint. For $(|\lambda_{13}|,|\lambda_{23}|)$, most of the points lie in the right-top region ranging from about 0.1 to 3.14 for most cases such as Case I , Case II and Case VI , where we have similar shapes of the parameter space no matter which couplings are negative. For other Cases, the choice of different couplings to be negative will make a difference in the parameter space. Particularly, for $\lambda_{13},\lambda_{23}\leqslant 0$ of Case IV, $\lambda_{12},\lambda_{13},\lambda_{23}\leqslant 0$ of Case V, $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$ of Case VII, $\lambda_{13},\lambda_{23},\lambda_{14},\lambda_{24}\leqslant 0$ of Case VIII, the larger $|\lambda_{13}|$ and $|\lambda_{23}|$ are excluded. In other words, the parameter space is more constrained as long as $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$ for the copositive criteria. The reason is that the processes $S_{i}S_{i}\to h_{i}h_{i}(i=1,2)$ are efficiently influenced by the interference effect in the case of $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$, and dark matter relic density will be under-abundant due to the larger annihilation processes arising from the larger couplings. For $(|\lambda_{14}|,|\lambda_{24}|)$, the interference effect on the cross section is less efficient compared with the for the low Higgs mass and SM vev, and the shapes of the parameter space of $(|\lambda_{14}|,|\lambda_{24}|)$ are almost similar among the different cases according to Fig. 2 to Fig. 10. We focus on the copositive criteria on the quartic couplings and do not show the viable parameter space of the masses of the scalars since for the TeV scale dark matter mass, the viable parameter space satisfying dark matter constrained is more flexible.
6 Summary and Outlook
@@
Weakly interacting massive particles as dark matter are facing serious challenges that no signals have been found in the direct detection experiments so far, and one solution to such a problem is multi-component dark matter models, which include two or more dark matter species. On the other hand, the multi-component dark matter should be constrained by theoretical constraints and experiment constraints. Particularly, for the extended scalar dark matter models. the vacuum stability demand the scalar potential has to be bounded from below, which can put stringent constraints on the parameter space. What’s more, the copositive criteria allow to the derivation analytic necessary and sufficient vacuum stability conditions for the couplings, and different copositivity criteria will contribute to different viable parameter spaces depending on the choice of the signs of the couplings.
@@
We consider a two-component scalar dark matter where dark matter particles are stabilized by extra $Z_{2}\times Z_{2}^{\prime}$ in this work. To obtain a stable vacuum, the $4\times 4$ matrix of the quadratic couplings should satisfy copositivity criteria. We pick up 17 different cases and focus on the quartic couplings in the model from the point of copositive criteria. The signs of the parameters will not only demand different copositivity criteria but can also influence the results of the cross section of the $2\to 2$ processes so that the dark matter relic density is due to the interference effect. We randomly scan the chosen parameter space with the copositivity criteria, dark matter relic density constraint and direct detection constraint. Among the 17 different cases, the shapes of the viable parameter space $|\lambda_{14}-|\lambda_{24}|$ are almost similar while for $|\lambda_{13}|-|\lambda_{23}|$ we can have different results due to the large interference effect of the $2\to 2$ processes of dark matter annihilation into scalars. As long as $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$, the large value of $|\lambda_{13}|$ and $|\lambda_{23}|$ are excluded for the too large cross section so that dark matter density being under-abundant.
@@
As we can see in this work, the different choices of the signs of couplings can indeed contribute to different parameter spaces with copositivity criteria. Our model is though simple while the complete discussion about the copositive criteria is complex for the different permutations of the indexes. For a more complex model, copositivity allows us to find analytic necessary and sufficient vacuum stability conditions so that estimate the parameter space theoretically.
@@ Acknowledgements.
Hao Sun is supported by the National Natural Science Foundation of China (Grant No. 12075043, No. 12147205).
@@ A.3 Renormalization Group Equations of the quartic couplings
In this part, we give the renormalization Group Equations (RGEs) of the model with SARAHStaub:2015kfa at one-loop level, and the beta functions of the quartic couplings are given by:
$\beta_{\lambda_{24}}$ $=+2\lambda_{12}\lambda_{14}-\frac{9}{10}g_{1}^{2}\lambda_{24}-\frac{9}{2}g_{2}^{2}\lambda_{24}+8\lambda_{22}\lambda_{24}+4\lambda_{24}^{2}+2\lambda_{23}\lambda_{34}+12\lambda_{24}\lambda_{44}+6\lambda_{24}y_{t}^{2}$ (39)
$\beta_{\lambda_{14}}$ $=-\frac{9}{10}g_{1}^{2}\lambda_{14}-\frac{9}{2}g_{2}^{2}\lambda_{14}+8\lambda_{11}\lambda_{14}+4\lambda_{14}^{2}+2\lambda_{12}\lambda_{24}+2\lambda_{13}\lambda_{34}+12\lambda_{14}\lambda_{44}+6\lambda_{14}y_{t}^{2}$ (40)
$\beta_{\lambda_{12}}$ $=2\lambda_{13}\lambda_{23}+4\lambda_{12}^{2}+4\lambda_{14}\lambda_{24}+8\lambda_{11}\lambda_{12}+8\lambda_{12}\lambda_{22}$ (41)
$\beta_{\lambda_{23}}$ $=2\lambda_{12}\lambda_{13}+4\Big{(}2\lambda_{22}\lambda_{23}+2\lambda_{23}\lambda_{33}+\lambda_{24}\lambda_{34}+\lambda_{23}^{2}\Big{)}$ (42)
$\beta_{\lambda_{13}}$ $=2\lambda_{12}\lambda_{23}+4\lambda_{13}^{2}+4\lambda_{14}\lambda_{34}+8\lambda_{11}\lambda_{13}+8\lambda_{13}\lambda_{33}$ (43)

### v4

@@ V.2 Theortical constraint of copositive criteria
The case (4) belongs to Case VIII with $\lambda_{12}\geqslant 0,\lambda_{13},\lambda_{14},\lambda_{23},\lambda_{24}\leqslant 0$, one can define:
$A_{11}=\lambda_{22}(\lambda_{33}\lambda_{12}^{2}-2\lambda_{13}\lambda_{23}\lambda_{12}+\lambda_{11}\lambda_{23}^{2}),A_{12}=\lambda_{22}(\lambda_{33}\lambda_{12}-\lambda_{13}\lambda_{23}),A_{13}=\lambda_{22}(\lambda_{34}\lambda_{12}-\lambda_{23}\lambda_{14}),$
$A_{22}=\lambda_{33}\lambda_{22}-\lambda_{23}^{2},A_{23}=\lambda_{22}\lambda_{34}-\lambda_{23}\lambda_{24},A_{33}=\lambda_{44}\lambda_{22}-\lambda_{24}^{2}.$
Therefore, the copositive criteria demands that:
$A_{11}\geqslant 0,A_{22}\geqslant 0,A_{33}\geqslant 0,A_{12}+\sqrt{A_{11}A_{22}}\geqslant 0,A_{13}+\sqrt{A_{11}A_{33}}\geqslant 0,A_{23}+\sqrt{A_{22}A_{33}}\geqslant 0,$ (27)
$A_{13}\sqrt{A_{22}}+A_{23}\sqrt{A_{11}}+\sqrt{2(A_{12}+\sqrt{A_{11}A_{22}})(A_{13}+\sqrt{A_{11}A_{33}})(A_{23}+\sqrt{A_{22}A_{33}})}+\sqrt{A_{11}A_{22}A_{33}}+A_{12}\sqrt{A_{33}}\geqslant 0.$ (28)
As we mentioned above, to obtain a stable vacuum, copositive criteria can put stringent constraints on the parameter space, and different copositive criteria will contribute to different viable parameter space which depends on the signs of the quartic couplings.
V.3 Results
In this part, we discuss the copositive criteria, dark matter relic density and direct detection constraint on the model. We consider the four different cases of the copositive criteria and focus on the viable parameter space of $(\lambda_{13},\lambda_{14})$. We fix $v_{1}=2$ TeV, $m_{h_{2}}=1.5$ TeV and $\sin\theta=0.01$ so that we can obtain $\lambda_{33}=0.2812$, $\lambda_{34}=0.0454$ and $\lambda_{44}=0.1309$. In addition, we set $|\lambda_{12}|=1$, $m_{2}=3.5$ TeV and fix $\lambda_{23}$ as well as $\lambda_{24}$ constants. Note that we have chosen the small $\sin\theta$ so that the contribution of $S_{i}S_{i}\to h_{1}h_{2}(i=1,2)$ is suppressed. The expression of the processes related to dark matter production in the limit of $\sin\theta\to 0$ can be found in App. A.4. We randomly scan the parameter space with
$m_{1}\subset\{1~\mathrm{TeV},2~\mathrm{TeV},3~\mathrm{TeV}\},|\lambda_{13}|\subset[0.0001,3.14],|\lambda_{14}|\subset[0.0001,3.14].$
In our scans, we consider the model compatible with the observed dark matter relic density if the value, as given by micrOMEGAs, lies between 0.119 and 0.121.
In Fig. 3, we give the results of case (1), where we set $\lambda_{12}=1$ and $\lambda_{13},\lambda_{14}\geqslant 0$. According to Fig. 3(a), we fix $\lambda_{23}=3.14,\lambda_{24}=0.0787$ while in Fig. 3(b) we set $\lambda_{23}=2.4,\lambda_{24}=0.0603$. Note that here the values of $\lambda_{24}$ is the upper bound obtained by dark matter direct detection constraint in the absence of $S_{1}$. The colored lines correspond to the viable parameter space of $(\lambda_{13},\lambda_{14})$ satisfying relic density constraint with $m_{1}=1,2,3$ TeV, and the respective dashed lines are the upper bound of $\lambda_{14}$ arising from direct detection constraint. The region of the colored lines below the representative dashed lines correspond to the parameter space both satisfying relic density and direct detection constraint. Direct detection result puts the most stringent constraint on the parameter space. Note that $\lambda_{13}$ is related to processes involving $h_{2}$ while $\lambda_{14}$ is related to processes involving SM Higgs ($h_{1}$). According to Fig. 3, when $\lambda_{13}$ is small, the value of $\lambda_{14}$ is approximately equal to a constant larger than 0.1 under relic density constraint, which means processes related to $h_{1}$ is dominant over dark matter production while the processes related to $h_{2}$ are less efficient. With the increase of $\lambda_{13}$, $\lambda_{14}$ decrease sharply as long as $\lambda_{13}$ approximate to 1, where processes related to $h_{2}$ is so efficient that $\lambda_{13}$ is constrained strictly and $\lambda_{14}$ has a lower bound respectively. Notice that direct detection constraint excludes the region that $\lambda_{13}$ is much smaller than 1, which means $h_{2}$ plays an important role in determining dark matter production within the viable parameter space. For the heavier $m_{1}$ with $m_{1}=2$ TeV and 3 TeV, the upper bound of $\lambda_{13}$ is larger to obtain the correct relic density as the lower bound of $\lambda_{14}$ is larger. However, in the case of $m_{1}=1$ TeV smaller than $m_{h_{2}}$, the process of $S_{1}S_{1}\to h_{2}h_{2}$ is highly suppressed and $\lambda_{13}$ should be much larger to meet relic density constraint. Respectively, the lower bound of $\lambda_{14}$ is larger according to Fig. 3(a). In Fig. 3(b) with $\lambda_{23}=2.4,\lambda_{24}=0.0603$, we have a similar conclusion, where the viable value of $\lambda_{13}$ is at $\mathcal{O}(1)$ level. However, the lower bound of $\lambda_{14}$ is much larger since the density of $S_{2}$ is larger in this case and one needs stronger interaction related to $S_{1}$ to obtain the correct relic density.
Figure 3: Results of case (1) with $\lambda_{12}=1,\lambda_{13},\lambda_{14}\geqslant 0$, where we fix $\lambda_{23}=3.14,\lambda_{24}=0.0787$ in the left picture, and $\lambda_{23}=2.4,\lambda_{24}=0.0603$ in the right picture. The colored lines correspond to the viable parameter space of $(\lambda_{13},\lambda_{14})$ satisfying relic density constraint with $m_{1}=1,2,3$ TeV, and the respective dashed lines are the upper bound of $\lambda_{14}$ arising from direct detection constraint.
Figure 4: Results of case (2) with $\lambda_{12}=1,\lambda_{13},\lambda_{14}\leqslant 0$, where we fix $\lambda_{23}=3.14,\lambda_{24}=0.0787$ in the left picture, and $\lambda_{23}=2.4,\lambda_{24}=0.0603$ in the right picture. The blue-shadowed region corresponds to the parameter space satisfying the copositive criteria of case ii. The colored lines correspond to the viable parameter space of $(|\lambda_{13}|,|\lambda_{14}|)$ satisfying relic density constraint with $m_{1}=1,2,3$ TeV, and the respective dashed lines are the upper bound of $|\lambda_{14}|$ arising from direct detection constraint.
Figure 5: Results of case (3) with $\lambda_{12}=-1,\lambda_{13},\lambda_{14}\leqslant 0$, where we fix $\lambda_{23}=3.14,\lambda_{24}=0.0787$ in the left picture, and $\lambda_{23}=2.4,\lambda_{24}=0.0603$ in the right picture. The blue-shadowed region corresponds to the parameter space satisfying the copositive criteria of case iii. The colored lines correspond to the viable parameter space of $(|\lambda_{13}|,|\lambda_{14}|)$ satisfying relic density constraint with $m_{1}=1,2,3$ TeV, and the respective dashed lines are the upper bound of $|\lambda_{14}|$ arising from direct detection constraint.
Figure 6: Results of case (4) with $\lambda_{12}=1,\lambda_{13},\lambda_{14},\lambda_{23},\lambda_{24}\leqslant 0$, where we fix $\lambda_{23}=-0.9,\lambda_{24}=-0.023$ in the left picture, and $\lambda_{23}=-0,75,\lambda_{24}=-0.194$ in the right picture. The blue-shadowed region corresponds to the parameter space satisfying the copositive criteria of case iii. The colored lines correspond to the viable parameter space of $(|\lambda_{13}|,|\lambda_{14}|)$ satisfying relic density constraint with $m_{1}=1,2,3$ TeV, and the respective dashed lines are the upper bound of $|\lambda_{14}|$ arising from direct detection constraint.
We give the results of case (2) in Fig. 4 with $\lambda_{12}=1,\lambda_{13},\lambda_{14}\leqslant 0$, where we fix $\lambda_{23}=3.14,\lambda_{24}=0.0787$ in the left picture, and $\lambda_{23}=2.4,\lambda_{24}=0.0603$ in the right picture. The colored lines correspond to the viable parameter space of $(|\lambda_{13}|,|\lambda_{14}|)$ satisfying relic density constraint with $m_{1}=1,2,3$ TeV, and the respective dashed lines are the upper bound of $|\lambda_{14}|$ arising from direct detection constraint. The blue-shadowed region corresponds to the parameter space satisfying copositive criteria. Since the cross-section of $S_{i}S_{i}\to h_{i}h_{i}(i=1,2)$ is determined by Higgs-mediated processes, the quartic term $|S_{i}|^{2}|h_{i}|^{2}$ and the t-channel processes exchange by dark matter, where the sign of the couplings $\lambda_{13}$ and $\lambda_{14}$ can make difference in the interference terms of these diagrams, therefore the viable parameter space will be different. According to Fig. 4(a), the smaller $|\lambda_{13}|$ is excluded for dark matter density being over-abundant due to the interference effect. Similarly, the viable $|\lambda_{13}|$ value is constrained within a narrow region while $|\lambda_{14}$ value is more flexible but sharply decreases with the increase of $|\lambda_{13}|$ by direct detection constraint as the results of case i, where $h_{2}$ plays a dominant role in determining dark matter production. On the other hand, for the heavier $m_{1}$, the upper bound of $|\lambda_{13}|$ is much smaller to obtain the correct relic density compared with the results of case i in Fig. 3(a) due to the interference effect. What’s more, copositive criteria in case ii can constrain the parameter space strictly and for $m_{1}=1$ TeV, the allowed $(|\lambda_{13}|,|\lambda_{14}|)$ is almost excluded as we can see in Fig. 4(a). For Fig. 4(b) with $\lambda_{23}=2.4,\lambda_{24}=0.0603$, the copositive criterial naturally excludes small $|\lambda_{13}|$ from the point of theoretical constraint in the case of $m_{1}=3$ TeV. Moreover, for $m_{1}=1$ TeV, the allowed parameter space of $(|\lambda_{13}|,|\lambda_{14}|)$ are completely excluded by the copositive criterial.
In Fig. 5 (a)and Fig. 5(b), we give the results of case (3) with $\lambda_{12}=-1,\lambda_{13},\lambda_{14}\leqslant 0$, where we fix $\lambda_{23}=3.14,\lambda_{24}=0.0787$ in the left picture, and $\lambda_{23}=2.4,\lambda_{24}=0.0603$ in the right picture. The difference between case (3) and case (2) is $\lambda_{12}$ set to be negative, the combined contribution of the Higgs-mediated processes and the quartic term $S_{1}^{2}S_{2}^{2}$ will determine the cross-section of $S_{1}S_{1}\to S_{2}S_{2}$, and the sign of $\lambda_{12}$ will affect the interference terms. However, such interactions between $S_{1}$ and $S_{2}$ just adjust he relative relic density of $S_{1}$ and $S_{2}$, and make little difference in the viable parameter space of $\lambda_{13}$ and $\lambda_{14}$. Therefore, we almost have similar results with case (2) as we can see in Fig. 5 (a)and Fig. 5(b).
We give the results of case (4) in Fig. 6 with $\lambda_{12}=1,\lambda_{13},\lambda_{14},\lambda_{23},\lambda_{24}\leqslant 0$. According to the copositive criteria of Case VIII, the value of $\lambda_{23}$ and $\lambda_{24}$ are also constrained. Within the chosen parameter space, one can obtain $|\lambda_{23}|\leqslant 0.94$ and $|\lambda_{24}|\leqslant 0.64$ with Eq. 27. In Fig. 6(a), we fix $\lambda_{23}=-0.9$ and $\lambda_{24}=-0.023$ and the allowed values of $|\lambda_{13}|$ as well as $|\lambda_{14}|$ are completely excluded by direct detection constraint and copositive criteria within the chosen parameter space in the case of $m_{1}=1$ TeV. On the other hand, for $m_{1}=3$ TeV, the allowed value of $|\lambda_{14}|$ is more flexible due to the interference effect. We set $\lambda_{23}=-0.75$ and $\lambda_{24}=-0.194$ in Fig. 6(b), where copositive criterial not just excludes the parameter space of $(|\lambda_{13}|,|\lambda_{14}|)$ in the case of $m_{1}=1$ TeV but also excludes the case of $m_{1}=3$ TeV regardless of direct detection constraint.
In conclusion, according to Fig. 3 to Fig. 6, we show the effect of copositive criteria on the parameter space. Different choices of the signs of the couplings will contribute to different copositive conditions, which can induce different viable parameter spaces due to the interference effect. We consider 4 different cases and randomly scan the chosen parameter space under dark matter relic density and direct detection constraints. Direct detection constraint puts the most stringent limit on the parameter space, which demands new Higgs $h_{2}$ plays a dominant role in determining dark matter relic density. On the other hand, copositive criterial can also constrain the allowed parameters space that the cases of $m_{1}=1$ TeV are excluded for different copositive conditions when some of the couplings are negative. Particularly, when we consider $\lambda_{12}=1,\lambda_{13},\lambda_{14}\leqslant 0,\lambda_{23}=-0.75$ and $\lambda_{24}=-0.194$, copositive criteria is more stringent on the viable parameters and the parameter space of $m_{1}=3$ TeV is completely excluded regardless of direct detection constraint. The high values of the couplings $\sim 1$ can imply a low energy validity for the model with two stable scalars, which will set limitations for the model at high energies. In the appendix, we perform a random scan and the viable parameter spaces with four different cases satisfying copositive criterial, relic density and direct detection constraint are shown. According to the results, the allowed values for the couplings can be at $\mathcal{O}(0.1)$ level, which are okay with high scale validity of the model.
VI Summary and Outlook
@@
Weakly interacting massive particles as dark matter are facing serious challenges that no signals have been found in the direct detection experiments so far, and one solution to such a problem is multi-component dark matter models, which include two or more dark matter species. On the other hand, the multi-component dark matter should be constrained by theoretical constraints and experiment constraints. Particularly, for the extended scalar dark matter models. the vacuum stability demand the scalar potential has to be bounded from below, which can put stringent constraints on the parameter space. What’s more, the copositive criteria allow the derivation analytic necessary and sufficient vacuum stability conditions for the couplings, and different copositivity criteria will contribute to different viable parameter spaces depending on the choice of the signs of the couplings.
@@
We consider a two-component scalar dark matter model in this work. To obtain a stable vacuum, the $4\times 4$ matrix of the quadratic couplings should satisfy copositivity criteria. Provided that direct detection results put stringent constraints on the parameter space, discussion about the related parameters can indeed be simplified greatly and we can just focus on 4 different cases related to copositive criteria depending on the signs of the quadratic couplings in the model. The signs of these parameters will not only demand different copositivity criteria but can also influence the results of the cross-section of the $2\to 2$ processes and induce different viable parameter space due to the interference effect. We randomly scan the chosen parameter space with dark matter relic density and direct detection constraints. For the 4 different cases, the allowed values of $(|\lambda_{13}|,|\lambda_{14}|)$ are different but all demand the new Higgs $h_{2}$ play a dominant role in determining dark matter relic density within the viable parameter space. In addition, we perform a random scan in the appendix under dark matter relic density and direct detection constraint, there are a few points ”survived” when copositive criterial is considered, and for different criterials, the allowed values for the couplings are different according to the results. In other word, the copositive criteria can also put stringent constraints on the parameter space as long as some quartic coupling is negative.
@@
As we can see in this work, different choices of the signs of couplings can indeed contribute to different parameter spaces with copositive criteria. The model in this work is simple but the discussion about copositive criterial is complex for the different permutations of the indexes. For a more complex model, copositivity allows us to find analytic necessary and sufficient vacuum stability conditions so that estimate the parameter space theoretically.
@@ Acknowledgements.
Hao Sun is supported by the National Natural Science Foundation of China (Grant No. 12075043, No. 12147205). XinXin Qi is supported by the National Natural Science Foundation of China (Grant No.12447162).
@@ A.3 Renormalization Group Equations of the quartic couplings
In this part, we give the renormalization Group Equations (RGEs) of the model with SARAHStaub (2015) at one-loop level, and the beta functions of the quartic couplings are given by:
$\beta_{\lambda_{24}}$ $=+2\lambda_{12}\lambda_{14}-\frac{9}{10}g_{1}^{2}\lambda_{24}-\frac{9}{2}g_{2}^{2}\lambda_{24}+8\lambda_{22}\lambda_{24}+4\lambda_{24}^{2}+2\lambda_{23}\lambda_{34}+12\lambda_{24}\lambda_{44}+6\lambda_{24}y_{t}^{2},$ (30)
$\beta_{\lambda_{14}}$ $=-\frac{9}{10}g_{1}^{2}\lambda_{14}-\frac{9}{2}g_{2}^{2}\lambda_{14}+8\lambda_{11}\lambda_{14}+4\lambda_{14}^{2}+2\lambda_{12}\lambda_{24}+2\lambda_{13}\lambda_{34}+12\lambda_{14}\lambda_{44}+6\lambda_{14}y_{t}^{2},$ (31)
$\beta_{\lambda_{12}}$ $=2\lambda_{13}\lambda_{23}+4\lambda_{12}^{2}+4\lambda_{14}\lambda_{24}+8\lambda_{11}\lambda_{12}+8\lambda_{12}\lambda_{22},$ (32)
$\beta_{\lambda_{23}}$ $=2\lambda_{12}\lambda_{13}+4\Big(2\lambda_{22}\lambda_{23}+2\lambda_{23}\lambda_{33}+\lambda_{24}\lambda_{34}+\lambda_{23}^{2}\Big),$ (33)
$\beta_{\lambda_{13}}$ $=2\lambda_{12}\lambda_{23}+4\lambda_{13}^{2}+4\lambda_{14}\lambda_{34}+8\lambda_{11}\lambda_{13}+8\lambda_{13}\lambda_{33}.$ (34)